OpenStudy (anonymous):

Find the integral of dx/(sqrt(1+x)) from 0 to 8.

OpenStudy (anonymous):

let u = 1+x du = dx ∫du / sqrt(u) = ∫ u^(-1/2) du = 2u^(1/2) = 2sqrt(x+1) I'll let you do the computation

OpenStudy (anonymous):

How did you get 2u^(1/2)?

OpenStudy (anonymous):

Actually, let's just make u = x + 1, so du = dx, which means: du/sqrt(u), which is the same as: u^(-1/2) * du, now integrating: 2sqrt(u) Replacing u: 2sqrt(x + 1). Now if they want it from 0 to 8, we just solve as follows: 2sqrt(8 + 1) - 2sqrt(0 + 1) = 6 - 2 = ?

OpenStudy (anonymous):

If you need more help go here: http://www.youtube.com/watch?v=ZJ6sjCub4jg That video helps with that specific type of question :-)

OpenStudy (anonymous):

Thank you guys so much.

OpenStudy (anonymous):

Thank you for letting me help you! Signed: \[Mangorox\]